Solve the sytem:
2x-y=3
3x+3y=5
Solving for x and y , you get a fraction . I dont know how to do this problem , please help me .

Answer by lwsshak3(11628)

Solve the sytem:
2x-y=3
3x+3y=5
Solving for x and y
***
Use elimination method:
multiply 1st equation by 3 to eliminate y
6x-3y=9
3x+3y=5
add to eliminate y
9x=14
x=14/9
from 1st equation:
y=2x-3=28/9-3=28/9-27/9=1/9
..
x=14/9
y=1/9